# Dynamic Neural Curiosity Enhances Learning Flexibility for Autonomous Goal Discovery

### v1

@@ II-D Habituation and Persistence
In order to limit the time spent on exploration, our approach introduces a cognitive process called habituation. Indeed, there is no consensus as to how the LC switch from a tonic mode with high arousal and engages exploration to a phasic mode (exploitation). Here, we propose to model that switch by an habituation paradigm where the phasic activation from LC happens only when the robot’s perception of an object does not produce a novelty effect anymore. Habituation is well studied by researchers, especially in developmental psychology by evaluating how long infants examine a new stimuli. Research first focused on the link between familiarity of the perception and exploratory behavior [57] and concluded that infants that have been familiarized to some objects prefer to look at new ones. A similar study [58] confirms that infants of seven and twelve months old have a decrease of attention when the objects become familiar. Interestingly, the study determined that the looking duration was not dependent on the age but more on the familiarity of the object. If these findings concern an exploration behavior and how habituation can influence it, the duration at which an individual learns a new skill must be taken into account.
@@
Cognitive persistence is a fundamental aptitude if one wants to learn a new goal. Observing how individuals persevere to learn something new can help to decide how and if a learning is actually fruitful. Regarding developmental psychology, the measure of persistence is investigated by assigning tasks to infants and evaluating how long they engage in a learning activity [59]. Here, researchers established that there is a link between an infant’s persistence and competence. This means that infants who were more stimulated at home persist more toward goal oriented actions. On the opposite, infants with cognitive delay demonstrated less persistence, as well as engaging toward less challenging learning activities [60]. From a neuroscience perspective, there is little known about cognitive persistence. Teubner-Rhodes [61] investigated this ability to overcome task difficulty in adults. With the use of fMRI, she located persistence in the prefrontal cortex and concluded that adults with high perseverance demonstrated better performance and adaptive control for the task. Later, a study focused on persistence for multilingual people [62] and located this process in the inferior frontal gyrus and dorsal anterior cingulate cortex (dACC). This could suggest that in addition to dealing with uncertainty and prediction-error, ACC is also involved in cognitive persistence.
These cognitive processes are necessary for our approach in order to support the shift from tonic to phasic activation for habituation, and to determine how persistent the robot has to be while learning a skill. Each aspect of cognition described in this section will be modelled with the Dynamic Field Theory.
@@ II-E Dynamic Field Theory
In this work, we model the different cognitive process with neural dynamics and more especially Dynamic Field Theory (DFT) [14]. By doing so, the system exhibits a range of different behaviors by only tuning the intrinsic parameters of the neural fields. The theoretical framework mathematically models the evolution and activity in time of large populations of neurons. The approach has been successful to model complex cognitive processed, such as visual working memory [63], intentionality [64] or motor habituation [65]. The formulation and implementation of DFT is provided in the Appendix -A. Regarding the modelling of motions, we adopted the dynamic movement primitive framework.
@@ II-F Dynamic Movement Primitive
In this work, we control the motion of a robotic arm by using the Dynamic Movement Primitive framework (DMP) [66], as they can exhibit a wide range of motions with a limited number of parameters [67]. The approach uses a set of differential equations to represent a movement with the advantage of being robust against perturbation. More precisely, we adopt the formalism introduced by Pastor [68], which allows to reach the generalization of a motion by adapting a start and goal parameter. More details about the method and its implementation can be found in the appendix -B.
@@ III-A Overview
Figure 1: (Left) Architecture flow from goal discovery with bottom-up attention (new event) to learning with curiosity. (Right) Connectivity flow between all components in terms of exploration (goal discovery) and exploitation (goal learning).
The Locus Coeruleus (LC) architecture is the main component of our model. In the brain, this nucleus is directly responsible for attention and task commitment. Here, we link intrinsic motivation and the LC activity through the engagement in learning new goals with different types of attention. We propose to articulate and describe how these cognitive processes might work together in this section.
The LC exhibits two modes of neural activation : tonic for exogenous attention and a phasic spiking mode for top-down attention during task engagement. Because of the time scale abstraction level, the DFT framework can not precisely differentiate them as well as biological neurons. Then, we define a tonic activation as a sustained and uninterrupted excitation activity while the phasic mode exhibits a more sparse and localized activation in time. Figure 1-left depicts the architecture flow from goal discovery with bottom-up attention to learning with curiosity, followed by the robot. At first, the robot detects any salient stimuli that can take place in the camera field of view. Then, the process of goal discovery begins by performing action babbling around the object. During this stage, we can consider that the type of attention is exogenous, since the robot is sensitive to any stimulis (objects) in the environment. At the same time, an habituation mechanism reduces the sensitivity towards the objects, thus reducing the neural activity within the LC. Then, the robot selects a goal to learn with the help of the curiosity and persistence mechanisms. At that step, there is a shift from bottom-up to top-down attention since the model only focuses on observing outcomes from a particular object. Finally, the robot can switch attention to a new object and starts over the process of goal discovery. However, this last step could be ignored and the robot will choose to discover new goals on a previously seen object.
In this work, we define exploration as the process of goal discovery and exploitation as the stage to learn these goals (See Fig. 1-right). The perception and action module are involved in both steps by discovering new goals and processing already seen stimulis for the former, and generating or using a DMP (dynamic motion primitive) for the latter [66]. The perception system observes the outcome of an action by coupling the color of the object along with its motion in space. Then, the habituation module is in charge to determine if the observed stimulis is seen as new or if the robot is already habituated to this specific object. The persistence, error and learning progress module are taking place during exploitation. Persistence determines for how long the robot should attempt the learning of the current goal. The error module monitors the discrepancy between the output from the forward model and the actual perceptual outcome, hence supporting the neural computation of the learning progress. During both stages, the perception engine sends direct inputs to the forward and inverse models present in the error component. The following section presents details the Locus-Coeruleus model.
@@ III-B Locus Coeruleus
Figure 2: Overview of the Locus Coeruleus component driving goal discovery and learning.
@@
The LC mechanism is coordinating the engagement level regarding the discovery or the learning of goals (See Fig. 2). At any time, the active node always exhibits a supra threshold level of activation and excites the slow boost component, a particular form of memory trace that evolves depending on two nodes inputs (Appendix -A). This memory trace builds activation when the active node projects excitation, maintains the current level if both connected nodes are down, and decays the activation if only the threshold node is up. At the beginning, the slow boost module slowly rises the resting level of the phasic and tonic neural field (NF) and stops when a peak of activation appears, triggering the stop node. The tonic activity of neurons is depicted with the tonic NF and the excitation incoming from the habituation mechanism (scene and object selection NF). This neural field exhibits a supra-threshold peak of activation as long as the robot is not accustomed to an object. The phasic NF gathers activations from the error and learning progress models but with a higher resting level. This means that after discovering goals, the phasic NF will be the first to see the emergence of supra-threshold activations. Tonic NF has a resting level of -2 and phasic NF -1.05. This implies that the LC can see a peak in tonic NF when the cumulated learning progress and error are below 0.05. This indicates that the LC can drive into exploration again when all the skills are learned. After discovering several goals through bottom-up attention and being habituated to an object, the robot does not exhibit any learning progress. Then, it is necessary to bind the errors memory trace to the phasic NF in order to bootstrap the learning of previously discovered goals. We empirically chose the gain factor so the robot can engage in learning a skill if there is no learning progress ($\approx$ 30 percents of the errors present in the error module). However, we keep the errors contribution low enough to favor the learning of a task that has a small learning progress. By doing so, the robot can begin the learning of goals that do not exhibit any learning progress instead of performing several stage of exploitation for each goal beforehand to build up the different learning progresses. The one dimensional object decision NF represents the object’s color in which the robot should explore or engage in learning.
@@
By analyzing the connectivity flow of the model, we can see that we prioritize exploration the same manner that exogenous attention seems to be operating within the LC. As mentioned in the literature (II), the engagement in a task can only operate if there are no salient stimuli and thus a low level of bottom-up attention. The priority to exploration comes along with the tonic activation. Our approach demonstrates a sustained activation (i.e tonic) while discovering new goals, thus preventing any possibility to learn a specific skill. During exploitation, the model selects a goal to learn on a regular basis. This process stops any activation within the phasic NF and raises again the resting level to select the goal with the highest learning progress (or error in the absence of any learning progress). In this model, the persistence mechanism directly influence the goal selection and depict a phasic activity. Finally, if the robot is habituated and has learned all the goals discovered, the tonic NF will select a know object to perform exploration again (Appendix -C).
@@ III-C Perception and habituation
The perception module processes the object’s motion direction after an action. A perception is formalized by a 2D neural field, where the horizontal and vertical dimensions represent the object color and motion’s angle as a goal, respectively (see Appendix -E for more details).
Figure 3: Architecture of the perception mechanism. It gathers the goal detector and the Hebbian learning between a goal and a DMP.
The motion of the object is defined by an angle between a static reference vector and the observed vector’s motion of the object. The angle is bound between -$\pi$ and $\pi$ radians and scaled between [0,100] to fit the neural field feature space. It is important to emphasize that the detection of the object’s motion takes place in real-time, i.e., while the robot performs the action), delivering a sustained input to the goal perceived NF (Neural Field). In Fig. 3, an inhibition of return is taking place with goal perceived NF, WM goals and the goals NF. This avoid perceiving a stimuli as new when the goal has already been discovered. During exploration and if the stimuli is new, a dynamic motion primitive (DMP) is generated by the robot controller and activates the associated DMP neuron through a one-to-one connection. The Hebbian learning scheme associates the generated DMP with the perceived outcome (goal focus NF) with a reward-gated rule similar to [64] :
@@
The learning occurs when there are activations from DMP neuron, learning signal and the goal focus NF. They are respectively denoted by $\sigma(dmp_{i})$, $\sigma(lc:explore)$ and $\sigma(g(col,ang))$. The term $\sigma(g(col,ang))$ represents the sigmoid activation of the goal focus NF at the object color ($col$) and angle ($ang$) location. When the robot engages in learning a task (i.e., exploitation) the goals NF is inhibited so a peak of activation can emerge within the goal focus neural field. Then, the activation spreads through the synaptic weights and triggers the DMP neuron corresponding to a particular goal.
Figure 4: Neural flow of the habituation process decreasing the objects novelty in time.
@@
The habituation mechanism is tightly linked with the perception module and is inspired by the work of Perone and Spencer [69]. The mechanism demonstrates how infants become familiar to similar stimuli in time and lead to a decrease of novelty. Here, the stimuli are the color and the motion features (i.e., within the perception module). The former processes input given by the latter and drives the LC from a tonic to a phasic mode. The color of the object serves as feature within the object selection NF (Fig. 4). The new perception field receives activation from the goal color dimension as well as from the explore node inside the LC module, supporting a supra-threshold activation only during exploration. The condition of satisfaction node (CoS EOA) is only active for a brief period of time when the robot has finished an action. This node correspond to the term a(t) in Eqn. (4). When a new goal occurs in the scene, the visual memory slowly decays at the color feature location, diminishing the inhibition to the object selection field. On the contrary, if an action results without the perception of a new goal, the visual memory builds up an activation that will inhibit the object selection. In practice, this implies that the robot keeps the exploration of new goals as long as the inhibition is weak enough within the object selection NF. The main difference with the habituation model of Perone and Spencer rest in the coupling between WM colors and the visual memory trace. Here, the inhibition of the visual memory trace is directly linked to the WM colors while their model associate an inhibitory layer with an Hebbian mechanism to the object selection NF. In our model, it results with the object remaining habituated even if the object is out of sight (supra threshold activation within WM colors) and with the possibility to decay inhibition when a new stimuli is observed.
@@
Here, the use of dynamic neural fields provides the possibility to design a more robust perception system. The inhibition of return prevents to identify a previously discovered goal as a new one if it is seen again during exploration. By tuning the neural field parameters (i.e the standard deviation of peaks), it is possible to adjust the likelihood of two goals. With a high standard deviation, two similar goals have a lower probability to be distinguished from each other. In addition, dynamic field theory demonstrated excellent results in modelling brain processes as habituation. This is to our knowledge the first curiosity driven architecture that integrates such mechanism to favor exploration toward novel objects. The following section will describe the action and the persistence system.
III-D Action Formation
Figure 5: Overview of the action module. The model employs an inhibition of return mechanism to generate two poses during goal discovery. Only a single pose is produced and used with a DMP while learning.
@@ III-D Action Formation
The role of this process is to perform action babbling around the object located in the scene (Fig. 5). During exploration, the module generates two poses ($x_{1}$,$y_{1}$) and ($x_{goal}$,$y_{goal}$) that the end-effector will go through from a resting position ($x_{start}$,$y_{start}$). In this work, the z dimension is fixed at one centimeter above the table for a generated pose. The robot records the end-effector position from the resting pose until reaching the second pose and creates a DMP if it leads to the discovery of a new stimuli. While learning a skill (exploitation), the robot activates the DMP parameters of a goal and only needs to create a single pose.
Related to the neural fields flow, a peak of activation emerges within the position object NF when an object is in the camera field of view. This neural field along with the other two dimensional fields present in the module are representing the x and y position of the object in the robot space. Then, we apply a specific convolution between the position object output and the action formation NF. The result of this convolution serves as inhibition to create poses necessary for the DMP creation/activation. The input to the action formation NF can be seen as a reversed Mexican-hat shape centered at the object location (see Fig. 6.
@@
Figure 6: (Left) Mexican-hat convolution with inhibition applied to the object (input to action formation NF). (Middle) Generated pose after rising the resting level within the action formation NF. (Right) The input peak position provided to the robot controller (filtering NF).
@@
In order to generate a pose, we use a slow boost component as seen in the LC module (see Section III-B). The slow boost builds up an activation if signal action is active, and the activation decays when the robot signals the end of an action (robot controller). The reversed Mexican-hat shape input to action formation favors peaks of activation around the object location. The signal action node receives inhibition when the robot is moving and when a peak emerges within action formation. The working memory keeps track of the activations in time and projects the inhibition back to action formation. Together, these two fields provide an inhibition of return effect, avoiding to select the same pose twice. Here, we provide a specific value for the global inhibition of the working memory (Appendix -A), so several supra-threshold activations can take place at the same time without over population. This means that only a limited number of activations can be self-sustained. If a new peak emerges, one of the ”old” activations dissipates. The robot controller is in charge of delivering excitation after two poses during exploration and after a single pose for exploitation. The working memory receives inhibition in case of a new stimuli (transient goal) and when the persistence mechanism reached the threshold.
Figure 7: General view of the persistence component. The module adapts the time spend on learning a goal depending on the error level of the forward model
@@
The persistence module bears a rhythm to the action module by signaling when to reset the inhibition of return while learning a specific goal. In the case where the goal error does not decrease, this module stops any further trial to learn a particular goal and clear the memory from all the generated pose. The core of the process lays with the threshold, force and action nodes. Threshold node is a classic zero dimensional neural field, transient action and force nodes are memory traces with a slow build up and fast decay for the former, and a quick build up and long decay for the latter. Threshold node receives inhibition from the local error level (i.e the error of the current goal) and excitation from the transient action. However, the current goal NF output is either zero or one because of the use of the absolute sigmoid as kernel activation. The output from the current goal is always higher than the local goal error who is bounded by the hyperbolic tangent function (section III-E). After each motion, the activation level within transient action gradually increases and spreads to the threshold node. Once the level of transient action overcomes the inhibition of the local error, it excites the attempts memory trace (a(t) term in Eqn. (4)) and leads to an update of the attempts memory trace. To resume, no activation happens within the attempts MT at the beginning of learning, meaning that there is no inhibition toward the LC:phasic. After several trials and if the error remains high, the inhibition will prevent the LC from pursuing the learning of that particular goal in the immediate future. However, the inhibition of a particular goal decreases after focusing on a different skill. By tuning the build up time ($\tau_{+}$) of transient action, it is possible to modulate the number of motion attempts. If $\tau_{+}$ is large, the robot will significantly increase the number of trial to learn a goal. By setting a high value for $\tau_{-}$ (attempts MT), the error level will decay slower. This would extend the time scale in which a goal can be selected again for a learning phase and a high goal error can expand it even further. Finally, the force node is triggered under the excitation of the threshold node and will inhibits the IOR and the activation node in LC to avoid rising the resting level of the slow boost component within the LC. This last inhibition essentially delay for a short time the next goal selection while the IOR activations within the action module are not completely down.
@@
Regarding the action module, one of the technical novelty rests in the use of the convolution applied to the object as well as the inhibition of return. This agency provides a simple and efficient way to generate poses around the object and provide a mean to focus the random exploration [36]. As for the persistence mechanism, a key novelty is to provide a mechanism that could adapt depending on the learning behavior. The use of dynamic neural fields support the flexibility to modulate the number of attempts made by the robot as well as the inhibition of the goal for future selection if the error remains significant. As a result, it is possible to observe diverse learning behavior by tuning the parameters of these neural fields (see sec. V-C)

### v2

@@ II-D Habituation and Persistence
In order to limit the time spent on exploration, our approach introduces a cognitive process called habituation. In fact, there is yet no evidence on how the MRN might influence the LC to switch from a tonic mode with high arousal to a phasic mode (exploitation). Here, we propose to model this switch by a habituation paradigm where the phasic activation from LC happens only when the robot’s perception of an object does not produce a novelty effect anymore. Researchers, particularly in developmental psychology, study habituation by measuring the duration of infant attention to novel stimuli [63]. The research on the topic first focused on the link between perception familiarity and exploratory behavior [64] and concluded that infants who have become familiar with some objects prefer to look at new ones. A similar study [65] confirms that babies between the ages of seven and twelve months have a decrease in attention when objects become familiar. Interestingly, the study determined that the look duration did not depend on age but rather on the familiarity of the object. Habituation is an essential mechanism for cognitive function [66], and simulation of this mechanism supports the reproduction of developmental studies [67]. In this work, we will adopt and modify the model proposed by Perone and Spencer [68] to avoid the dishabituation of objects when they are out of sight. In addition, the proposed model will avoid reproducing the familiarity effect for this study, as it can complicate the study of the goal discovery and learning stages.
@@
Cognitive persistence is a fundamental aptitude if one wants to learn a new goal. Observing how individuals persevere to learn something new can help to decide how and whether learning is actually fruitful. Regarding developmental psychology, the measure of persistence is investigated by assigning tasks to infants and evaluating how long they engage in a learning activity [69]. Here, researchers established that there is a link between an infant’s persistence and competence. This means that infants who were more stimulated at home persist more in goal-oriented actions. In contrast, infants with cognitive delay demonstrated less persistence, as well as a desire to participate in less challenging learning activities [70]. From a neuroscience perspective, Teubner-Rhodes investigated this ability to overcome task difficulty in adults [71]. Using fMRI, she located persistence in the prefrontal cortex and concluded that adults with high perseverance demonstrated better performance and adaptive control for the task. Later, a study focused on persistence for multilingual people [72] and located this process in the inferior frontal gyrus and dorsal anterior cingulate cortex (dACC). This could suggest that, in addition to dealing with uncertainty and prediction error, ACC is also involved in cognitive persistence. The MRN is involved in cognitive persistence [62], and the suppression of inhibitory neurons leads to more perseverative behavior. More specifically, the serotonin neurotransmitter appears to have a strong influence on perseverative states through widespread inhibition [73, 74]. These studies question the roles of the MRN along with the LC as they seem to partially overlap regarding task engagement, but also seem complementary during learning.
In this work, we intend to propose a simple model of persistence (i.e., perseverance) to regulate the learning of a goal via the involvement of prediction errors. The purpose is to endow the robot with the ability to dynamically allocate time to learn a goal depending on learning progress and prediction error. Each aspect of cognition described in this section will be modeled with dynamic field theory.
@@ II-E Dynamic Field Theory
In this work, we model the different cognitive processes with neural dynamics and, more especially, Dynamic Field Theory (DFT) [16]. By doing so, the system exhibits a range of different behaviors by tuning only the intrinsic parameters of the neural fields. The theoretical framework mathematically models the evolution and activity in time of large populations of neurons. The approach has been successful in modeling complex cognitive functions, such as visual working memory [75], intentionality [76], or motor habituation [77]. The formulation and implementation of DFT are provided in the Appendix -A. Regarding the modeling of motions, we adopted the dynamic movement primitive framework.
@@ II-F Dynamic Movement Primitive
In this work, we control the motion of a robotic arm using the Dynamic Movement Primitive Framework (DMP) [78], as it can exhibit a wide range of motions with a limited number of parameters [79]. The approach uses a set of differential equations to represent a movement, with the advantage of being robust against perturbation. More precisely, we adopt the formalism introduced by Pastor [80], which allows the generalization of a motion by adapting a start and goal parameter. More details about the method and its implementation can be found in the Appendix -B.
@@ III-A Overview
Figure 1: Timecourse for the discovery and learning of goals.
The Locus Coeruleus (LC) architecture is an essential component of our model. In the brain, this nucleus is directly responsible for attention and task engagement. Here, we link intrinsic motivation and LC activity through engagement in learning new goals with different types of attention. In this section, we propose to articulate and describe how these cognitive processes might work together.
Figure 2: Overview of the general architecture with the cognitive processes. Connections between the mechanisms are detailed within each component.
The LC exhibits two modes of neural activation: tonic for exogenous attention and phasic spiking mode for top-down attention during task engagement. Due to the time-scale abstraction level, the DFT framework cannot precisely differentiate them as well as biological neurons. Then, we define a tonic activation as a sustained and uninterrupted excitation activity, while the phasic mode exhibits a more sparse and localized activation over time. Figure 1 depicts the architecture steps from goal discovery with bottom-up attention to learning with curiosity. At first, the robot detects objects with the camera and applies a color filter. This information is formalized as neural activity and projected to the habituation component that selects an object to begin exploration. Then, the goal discovery process begins by performing action babbling around the object. During this stage, we can consider that the type of attention is exogenous, since the robot is sensitive to any stimuli related to the selected object. At the same time, a habituation mechanism reduces the sensitivity towards the objects, thus reducing the neural activity within the LC. The robot then selects a goal to learn with the help of the curiosity and persistence mechanisms. At that step, there is a shift from bottom-up to top-down attention, since the model only focuses on observing outcomes from a particular object. Finally, the robot can switch attention to a new object and start over the process of goal discovery. However, this last step could be ignored, and the robot will choose to discover new goals on a previously seen object.
@@ III-B Locus Coeruleus
The LC mechanism coordinates the engagement in the discovery or learning of goals. At any time, the active node always exhibits a suprathreshold level of activation and excites the slow boost component (see Fig. 2), a particular form of memory trace that evolves depending on two node inputs (Appendix -A). This memory trace builds activation when the active node projects excitation, maintains the current level if both connected nodes are down, and decays the activation if only the threshold node is up. At first, the slow boost module slowly increases the resting level of the phasic and tonic neural field (NF) and stops when a peak of activation appears, triggering the stop node. The tonic activity of neurons is depicted with the tonic NF and the excitation incoming from the habituation mechanism (object selection NF). This neural field exhibits a peak of activation as long as the robot is not accustomed to an object. The phasic NF gathers activations from the error and learning progress modules, but with a higher resting level. This means that after discovering goals, the tonic NF will be the first to see the emergence of suprathreshold activations. Tonic NF has a resting level of -2 and phasic NF -1.05. This implies that LC can see a peak in tonic NF when cumulative learning progress and error are below 0.05. This indicates that the LC can go back to exploration when all the goals have been learned. After discovering several goals through bottom-up attention and being habituated to an object, the robot does not show any learning progress. Then, it is necessary to bind the error memory trace to the phasic NF to bootstrap the learning of previously discovered goals. We empirically chose the gain factor so that the robot can engage in learning a goal if there is no learning progress ($\approx$ 30 percents of the errors present in the error module). However, we keep the error contribution low enough to favor the learning of a task that has a small learning progress. By doing so, the robot can begin to learn goals that do not exhibit any learning progress, instead of performing several stages of exploitation for each goal beforehand to build up the different learning progresses. The one-dimensional object decision NF represents the object’s color in which the robot should explore or engage in learning.
@@
Here, we prioritize exploration in the same manner that exogenous attention seems to be operating within the LC. As mentioned in the literature (II), the engagement in a task can only operate if there are no salient stimuli and thus a low level of bottom-up attention. The priority of exploration comes with the tonic activation. Our approach demonstrates sustained activation (i.e., tonic) while discovering new goals, thus preventing the possibility to learn a specific goal. During exploitation, the model selects a goal to learn regularly. This process stops any activation within the phasic NF and raises again the resting level to select the goal with the highest learning progress (or error in the absence of any learning progress). In this model, the persistence mechanism directly influences the selection of the goal and depicts the phasic activity. Finally, if the robot is habituated and has learned all the discovered goals, the tonic NF will select a known object to conduct the exploration again (Appendix -C).
@@ III-C Perception and habituation
The perception module processes the color and motion of the object. The colors of objects in sight are formalized as inputs to a one-dimensional field and sent to the habituation mechanism, where neural dynamics can perform the selection of an object. A goal is formalized by the motion direction of an object. This motion is characterized by an angle between a reference vector and a vector representing the motion of the object along the dimensions $x$ and $y$ (see Appendix -E). Here, we voluntarily design the goal as simply as possible to study the underlying mechanism of goal discovery and learning. The goal space is then formalized by a 2D neural field, where the vertical and horizontal dimensions represent the object color and motion angle as a goal, respectively (Figure 2). It is important to emphasize that the detection of the object’s motion takes place in real time (i.e., while the robot performs the action), delivering a sustained input to the goal perceived NF (Neural Field). In Figure 2, an inhibition of return occurs with the goal perceived NF, the WM goals, and the NF goals to avoid perceiving stimuli as new when the goal has already been discovered. During exploration, and if the stimulus is new, the robot controller generates a dynamic motion primitive (DMP) and activates the associated DMP neuron through a one-to-one connection. The Hebbian learning model links the created DMP to the observed outcome (goal focus NF) using a reward-gated rule similar to [76]:
@@
Learning occurs when there are activations from the DMP neuron, the explore node, and the goal focus NF. They are respectively denoted by $\sigma(dmp_{i})$, $\sigma(lc:explore)$ and $\sigma(g(col,ang))$. The term $\sigma(g(col,ang))$ represents the sigmoid activation of the goal focus NF at the object color (colcol) and angle (angang) location. When the robot engages in learning a goal (i.e., exploitation), the goals NF is inhibited, so a peak of activation can emerge within the goal focus neural field. The activation then spreads through the synaptic weights and triggers the DMP neuron corresponding to a particular goal.
@@
The habituation mechanism is closely related to the perception module and is inspired by the work of Perone and Spencer [68]. The mechanism slowly habituates the robot to repeated stimuli, such as observing the same goal from the same object. The LC switches from a tonic mode to a phasic mode because of habituation. The color of the object serves as a feature within the object selection NF (Fig. 2). The new perception field receives activation from the goal color dimension, as well as from the explore node inside the LC module, supporting a suprathreshold activation only during exploration. The Condition of Satisfaction Node (CoS) is only active for a brief period of time when the robot has finished an action (End of Action). This node corresponds to the term a(t) in Eqn. (4). When a new goal occurs in the scene, visual memory slowly decays at the color location, reducing the inhibition of the object selection field. In contrast, if an action results without the perception of a new goal, visual memory builds up an activation that will inhibit object selection NF. In practice, this implies that the robot keeps the exploration of new goals as long as the inhibition is weak enough within the object selection NF. The main difference from the habituation model of Perone and Spencer rests in the coupling between the WM colors and the visual memory trace. Here, the inhibition of the visual memory trace is directly linked to the WM colors, while their model associates an inhibitory layer with a Hebbian mechanism to the object selection NF. In our model, it results in the object remaining habituated even if it is out of sight, and with the possibility of decaying inhibition when new stimuli are observed.
@@
Here, the use of dynamic neural fields provides the possibility of designing a more robust perception system. The inhibition of return prevents the robot from identifying a previously discovered goal as a new one if it is seen again during exploration. By tuning the neural field parameters (i.e., the standard deviation of peaks), it is possible to adjust the likelihood of two goals. With an elevated standard deviation, two similar goals have a lower probability of being distinguished from each other. In addition, dynamic field theory demonstrated excellent results in modeling brain processes as habituation. This is, to our knowledge, the first curiosity-driven architecture that integrates such a mechanism to favor exploration toward novel objects. The following section will describe the action and the persistence system.
III-D Action and Persistence
@@
The role of this process is to perform action babbling around the object located in the scene (Fig. 2). During exploration, the module generates two poses ($x_{1}$$y_{1}$) and ($x_{goal}$$y_{goal}$) through which the end effector will go from a resting position ($x_{start}$$y_{start}$). In this work, the z dimension is fixed at one centimeter above the table for a generated pose. The robot records the end effector position from the resting pose until reaching the second pose and creates a DMP if it leads to the discovery of a new stimulus. During goal learning, the robot activates the DMP parameters of a goal and only needs to generate a single pose.
Regarding object perception, a filter based on the selected object is applied (i.e., from the habituation component). After a transformation to align the image to the robot space coordinate system, the filtered image serves as input to the position object NF (Fig. 3-left). Then, we apply a specific convolution between the position object output and the action formation NF. The result of this convolution then serves as an inhibition to action formation NF for the action babbling behavior (see Fig. 3-center).
@@
Figure 3: (Left) Object input (cube) to position object NF in the action component. (Middle) Mexican-hat convolution with inhibition applied to the object (input to action formation NF). (Right) Generated pose after raising the resting level within the action formation NF.
@@
To generate a pose, we use a slow boost component as seen in the LC module (see Section III-B). The slow boost builds up an activation if signal action is active, and the activation decays when the robot signals the end of an action (robot controller). The reversed Mexican-hat shape input to action formation favors activation peaks around the location of the object (Fig. 3-right). The signal action node receives inhibition when the robot is moving and when a peak emerges within the action formation NF. The working memory keeps track of the activations in time and projects the inhibition back to action formation. Together, these two fields provide an inhibition of return effect, avoiding selecting the same pose twice. Here, we provide a specific value for the global inhibition of working memory (Appendix -A), so several suprathreshold activations can take place at the same time without overpopulation. This means that only a limited number of activations can be self-sustained. If a new peak emerges, one of the ”old” activations dissipates. The robot controller is in charge of delivering excitation after two poses during exploration and after a single pose for exploitation. Working memory is inhibited in the event of new stimuli (transient goal) and when the persistence mechanism reaches the threshold.
@@
The persistence module carries a rhythm to the action module by signaling when to reset the inhibition of return while learning a specific goal. In the case where the goal error does not decrease, this module stops any further trial to learn a particular goal and clear the memory of all the generated poses. The core of the process lies with the threshold, force, and action nodes. The threshold node is a classic zero-dimensional neural field, transient action and force nodes are memory traces with a slow build-up and fast decay for the former, and a quick build-up and long decay for the latter. The threshold node receives inhibition from the local error level (i.e., the error of the current goal) and excitation from the transient action. However, the current goal NF output is either zero or one because of the use of the absolute sigmoid as kernel activation. The output of the current goal is always higher than the local goal error that is bounded by the hyperbolic tangent function (Section III-E). After each movement, the level of activation within the transient action gradually increases and spreads to the threshold node. Once the level of transient action overcomes the inhibition of the local error, it excites the attempts memory trace (a(t) term in Eqn. (4)) and leads to its update. To resume, no activation occurs within the attempts MT at the beginning of learning, meaning that there is no inhibition toward the LC:phasic NF. After several trials and if the error remains high, inhibition will prevent the LC from pursuing the learning of that particular goal in the immediate future. However, the inhibition of a particular goal decreases after focusing on a different goal. By tuning the build-up time ($\tau_{+}$) of the transient action node, it is possible to modulate the number of motion attempts. If $\tau_{+}$ is large, the robot will significantly increase the number of trials to learn a goal. By setting a high value for $\tau_{-}$ (attempts MT), the error level will decay more slowly. This would extend the time scale in which a goal can be selected again. Finally, the force node is activated under excitation of the threshold node and will inhibit the IOR and the activation node in the LC to avoid raising the resting level of the slow boost component within the LC. This last inhibition essentially delays for a short time the next goal selection while the IOR activations within the action module are not completely down.
@@
Regarding the action module, one of the technical novelties rests in the use of the convolution applied to the object, as well as the inhibition of return. This agency provides a simple and efficient way to generate poses around the object and provides a method to focus the random exploration [41]. As for the persistence mechanism, a key novelty is to provide a mechanism that could adapt depending on the learning behavior. The build-up of activation within the attempts NF is similar to the habituation mechanism and can be seen as a habituation to prediction error, on the difference that the prediction error level continuously changes during the learning of a goal. The use of dynamic neural fields supports the flexibility to modulate the number of attempts made by the robot, as well as the inhibition of the goal for future selection if the error remains significant. As a result, it is possible to observe diverse learning behavior by tuning the parameters of these neural fields (see sec. V-C)